Toyota decided to develop a new hybrid car. The project costs 50 billion yen to complete. The company's managers expected that the new car would bring 70 billion yen in additional sales. Toyota has already invested 20 billion yen when it suddenly finds out that the introduction of a similar car by Honda is expected to reduce Toyota's additional sales to 27 billion yen. What is Toyota's sunk cost?
Answer #1

The project costs 50 billion yen to complete. The company's managers expected that the new car would bring 70 billion yen in additional sales. Toyota has already invested 20 billion yen when it suddenly finds out that the introduction of a similar car by Honda is expected to reduce Toyota's additional sales to 27 billion yen.

Toyota's additional cost = (70-50)/(27-20)

= 20/7 = 2.86 Billion